OpenStudy (anonymous):

Is there any shortcut to find partials if we have a quadratic factor in denominator??/

OpenStudy (primeralph):

Example?

OpenStudy (kainui):

Put the denominator in the numerator with a -1 exponent. Then you can use the product and chain rules. That's how I roll.

OpenStudy (anonymous):

\[(2x ^{2}+3)/(x ^{2}+4)(x ^{2}+6) \]

OpenStudy (primeralph):

Why partials then if you have one variable?

OpenStudy (kainui):

I hope you're talking about the partial with respect to y!

hartnn (hartnn):

that example should be easy, you just put x^2 = y and solve like simple partial fractions, i think you are asking for something like (x^2+2x+4) (x^2+3x+6) in the denominator ?

OpenStudy (anonymous):

Normally what we do is equate this expression with \[(Ax+B)/(x ^{2}+4) + (Cx+D)/(x ^{2}+6)\]

OpenStudy (primeralph):

Partial fractions?

OpenStudy (anonymous):

@hartnn what if we have a linear expression in numerator

hartnn (hartnn):

ok, like 2x+3 / ()()

OpenStudy (anonymous):

yes....linear in numerator and quadratic in denominator

hartnn (hartnn):

the best method i know of is to take LCM and compare the co-efficients

hartnn (hartnn):

2x+3 = (Ax+B)(x^2+6) +(Cx+D)(x^2+4)

hartnn (hartnn):

so, if you compare the co-efficient of x^3, we get A+C = 0

hartnn (hartnn):

of x^2, >>>> B+D = 0

hartnn (hartnn):

and so on.... got it ? if others have an shorter way, they're welcome to share :)

OpenStudy (anonymous):

ya..exactly i know this method....but this goes long...

hartnn (hartnn):

not that i know of any shorter way.....i'll wait for others' responses..

OpenStudy (anonymous):

ok